# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v2

@@ II-B Enabling Mappings and Information
    Energy requirement: The transformation requires minimum energy
@@
    $E(f)\geq k_{B}T\ln\left(\frac{|\mathcal{S}(X)|}{|\mathrm{im}(f)|}\right),$ (1)
    where $\mathrm{im}(f)=\{f(x):x\in\mathcal{S}(X)\}$ is the image of $f$.
    For surjective mappings, $|\mathrm{im}(f)|=|\mathcal{S}(Y)|$, so
    $E(f)\geq k_{B}T\ln\left(\frac{|\mathcal{S}(X)|}{|\mathcal{S}(Y)|}\right).$ (2)
Physical interpretation: This reflects Landauer’s principle—compressing from $|\mathcal{S}(X)|$ states to $|\mathrm{im}(f)|$ states erases $\log_{2}\left(|\mathcal{S}(X)|/|\mathrm{im}(f)|\right)$ bits of information, requiring at least $k_{B}T\ln\left(|\mathcal{S}(X)|/|\mathrm{im}(f)|\right)$ joules of energy to be dissipated as heat.
Remark 1 (Energy Budget Interpretation).
The energy $E$ in the following theorem represents the total thermodynamic cost of the mapping process. It accounts for:
  • Erasure cost (irreversible): Information compression from $\mathcal{S}(O,T_{0})$ to $\mathcal{S}(C,T_{r})$ requires energy dissipation as specified by Landauer’s principle (Definition 3).
  • Maintenance cost (barrier): Sustaining the carrier state $C$ against thermal fluctuations at temperature $T$ requires an energy barrier proportional to the information stored. Following the Brillouin-Szilard principle [brillouin1956science][bennett1982thermodynamics], maintaining $n$ bits of information requires an energy barrier of at least $n\cdot k_{B}T\ln 2$.
For the lower bounds derived in Theorem 1, we use the total energy budget $E$ as a conservative estimate. A more refined analysis separating erasure and maintenance costs would strengthen these bounds but is not necessary for the inequalities presented here.
Definition 4 (Information).
Information is a six-tuple $I=\langle O,T_{o},S(O,T_{o}),C,T_{r},S(C,T_{r})\rangle$ where:
  • $O\subseteq W\cup M$: ontology (objective or subjective reality)
  • $T_{o}\subseteq\mathbb{R}^{+}$: occurrence time
  • $S(O,T_{o})$: state set of ontology at time $T_{o}$
  • $C\subseteq W$: carrier (objective reality)
  • $T_{r}\subseteq\mathbb{R}^{+}$: reflection time (with $T_{o}\subseteq T_{r}$)
  • $S(C,T_{r})$: state set of carrier at time $T_{r}$
satisfying $I:S(O,T_{o})\rightrightarrows S(C,T_{r})$ is an enabling mapping[xu2024research].
Theorem 1 (Energy-Constrained Representation Capacity).
For any enabling mapping $\mathcal{I}:S(O,T_{o})\rightrightarrows S(C,T_{r})$ requiring energy $E$, the following hold:
@@
    (Carrier capacity lower bound)
    $\log_{2}|\mathcal{S}(C,T_{r})|\geq\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (3)
@@
    (Information preservation bound) Under any joint distribution $P_{O,C}$ consistent with the deterministic mapping $C=f(O)$,
    $I(O;C)\leq\frac{E}{k_{B}T\ln 2},$ (4)
    and consequently,
    $H(O\mid C)\geq H(O)-\frac{E}{k_{B}T\ln 2}.$ (5)
@@
    (Uniform case—optimal compression) If $\mathcal{S}(O,T_{0})$ is uniformly distributed and the mapping partitions states uniformly (i.e., all fibers have equal size), then equality holds in (3):
    $\log_{2}|\mathcal{S}(C,T_{r})|=\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (6)
    For non-uniform partitions, inequality (3) is strict.
Proof.
We prove each part in turn.
(1) Carrier capacity lower bound.
By Definition 3, for the surjective mapping $\mathcal{I}:S(O,T_{o})\rightrightarrows S(C,T_{r})$, the energy satisfies
$E\geq k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (7)
Rearranging gives
$\ln|\mathcal{S}(C,T_{r})|\geq\ln|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T}.$ (8)
Converting to base-2 logarithms (dividing both sides by $\ln 2$) yields
$\log_{2}|\mathcal{S}(C,T_{r})|\geq\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2},$ (9)
which establishes (3).
Physical intuition: Given energy budget $E$, the mapping can erase at most $E/(k_{B}T\ln 2)$ bits of information. Therefore, the carrier must retain at least $\log_{2}|\mathcal{S}(O,T_{0})|-E/(k_{B}T\ln 2)$ bits, which determines the minimum carrier capacity.
(2) Information preservation bound.
Since $C=\mathcal{I}(O)$ is a deterministic function, $H(C\mid O)=0$, and thus
$I(O;C)=H(C)-H(C\mid O)=H(C)\leq\log_{2}|\mathcal{S}(C,T_{r})|.$ (10)
From thermodynamics, the maximum amount of information that can be maintained against thermal noise with energy $E$ is $E/(k_{B}T\ln 2)$ bits. [^1] This follows from the Brillouin-Szilard principle: maintaining information against thermal fluctuations requires an energy barrier proportional to $k_{B}T$ per bit. See Bennett [bennett1982thermodynamics][bennett2003notes] for a detailed thermodynamic analysis of information storage and the fundamental limits imposed by thermal noise. Therefore,
$I(O;C)\leq\frac{E}{k_{B}T\ln 2},$ (11)
proving (4).
Using the identity $H(O\mid C)=H(O)-I(O;C)$, we immediately obtain
$H(O\mid C)\geq H(O)-\frac{E}{k_{B}T\ln 2},$ (12)
which is (5).
(3) Uniform case.
Assume $O$ is uniformly distributed over $\mathcal{S}(O,T_{0})$, so $H(O)=\log_{2}|\mathcal{S}(O,T_{0})|$.
If the mapping $\mathcal{I}$ partitions $\mathcal{S}(O,T_{0})$ into fibers of equal size, then
$|\mathcal{I}^{-1}(c)|=\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\quad\text{for all }c\in\mathcal{S}(C,T_{r}).$ (13)
Under uniform partitioning, $C$ is also uniformly distributed, so $H(C)=\log_{2}|\mathcal{S}(C,T_{r})|$. The information erased is exactly
$\Delta I=H(O)-H(C)=\log_{2}\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right)\text{ bits}.$ (14)
By Landauer’s principle, the energy consumed is
$E=k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (15)
Solving for $\log_{2}|\mathcal{S}(C,T_{r})|$ gives
$\log_{2}|\mathcal{S}(C,T_{r})|=\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2},$ (16)
which is the equality (6).
For non-uniform partitions, the distribution of $C$ becomes non-uniform with $H(C)<\log_{2}|\mathcal{S}(C,T_{r})|$. The actual information erased is
$\Delta I=H(O)-H(C)>\log_{2}\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right),$ (17)
requiring energy
$E>k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (18)
Consequently, the inequality (3) becomes strict:
$\log_{2}|\mathcal{S}(C,T_{r})|>\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (19)
Physical interpretation: Uniform partitions maximize entropy production efficiency—they erase the minimum amount of information for a given compression ratio $|\mathcal{S}(O)|/|\mathcal{S}(C)|$, making them thermodynamically optimal. Non-uniform partitions waste energy by erasing more information than necessary.
II-C Logical Structure and Atomic Decomposition
Definition 5 (Logical System).
A logical system $\mathcal{L}=(\mathcal{A},\text{WFF},\vdash,\models)$ consists of[qiu2025research][xu2025information]:
  • $\mathcal{A}$: finite set of atomic propositions
  • WFF: well-formed formulas built from $\mathcal{A}$ and connectives $\{\neg,\land,\lor,\to\}$
  • $\vdash$: proof system (e.g., natural deduction, sequent calculus)
  • $\models$: model-theoretic semantics
Definition 6 (Logical Closure).
The logical closure $\text{Cn}(S)$ of a formula set $S\subseteq\text{WFF}$ is:
$\text{Cn}(S)=\{\varphi\in\text{WFF}:S\vdash\varphi\}$ (20)
Definition 7 (Atomic Decomposition).
For state $s$ with interpretation $I(s)\subseteq\text{WFF}$, the atomic decomposition $\text{Atom}(s)$ is the unique minimal set satisfying:
@@ II-C Logical Structure and Atomic Decomposition
    $\text{Atom}(s)\subseteq\mathcal{A}$ (atoms only)

### v3

@@ III-A Derivation Entropy and Information Content
    Proof system structure: Selecting from $|\mathcal{R}|$ rules per step suggests $\lambda\gtrsim\ln|\mathcal{R}|$ nats, as a structural lower bound on the information cost of inference.
TABLE I: Analogy Between Thermodynamic and Information-Theoretic Costs
Domain Discrete Unit Continuous Unit Conversion Factor
Thermodynamics 1 bit erasure Energy (J) $k_{B}T\ln 2$
Information Physics 1 inference step Information (nats) $\ln 2$
Assumption 3 (Classical Thermodynamic Regime).
The knowledge system operates in the classical thermodynamic regime: (1) Inference steps are logically irreversible at thermal equilibrium; (2) Each step erases $O(1)$ bits; (3) Landauer erasure dominates other costs. Under this assumption, $\lambda=\ln 2$ nats.
Corollary 3 (Derivation Entropy in Classical Regime).
Under Assumption 3: $H_{\mathrm{derive}}(\phi|\psi)=\mathrm{Dd}(\phi|\psi)\cdot\ln 2$.
Lemma 1 (Derivation Encoding).
Let $\mathcal{P}$ be a proof system with rule set $R$ and maximum rule arity $k$. Any derivation $\pi$ of length $n$ from $m$ premises satisfies:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)),$ (15)
where $c_{L}=\max\{\lceil\log_{2}|R|\rceil,k\}$.
A derivation $\pi=((r_{1},\vec{p}_{1}),\ldots,(r_{n},\vec{p}_{n}))$ consists of $n$ steps. Each step $(r_{i},\vec{p}_{i})$ encodes:
  • Rule: $r_{i}\in R$ in $\lceil\log_{2}|R|\rceil$ bits.
  • Premises: $\vec{p}_{i}=(p_{i,1},\ldots,p_{i,k_{i}})$ with $k_{i}\leq k$ pointers, each in $\{1,\ldots,m+i-1\}$, requiring $\leq k\cdot\lceil\log_{2}(m+n)\rceil$ bits total.
Per-step encoding: $\lceil\log_{2}|R|\rceil+k\cdot\lceil\log_{2}(m+n)\rceil\leq c_{L}\cdot(1+\log_{2}(m+n))$.
Over $n$ steps: $n\cdot c_{L}\cdot(1+\log_{2}(m+n))$.
Adding $O(\log(m+n))$ header bits for $m,n$, and noting that $n\cdot c_{L}\leq n\cdot c_{L}\cdot\log(m+n)$ when $\log(m+n)\geq 1$:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)).\qed$
Definition 8 (Information-Rich Regime).
Let $S_{O}$ be a knowledge base and let $q$ be a query (logical formula) over $S_{O}$. We define the conditional algorithmic entropy of $q$ given $S_{O}$ as
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr),$ (16)
where $K(\cdot\mid\cdot)$ denotes conditional Kolmogorov complexity with respect to a fixed universal Turing machine and a computable encoding scheme $\langle\cdot\rangle$.
We say that $q$ lies in the information-rich regime (with respect to $S_{O}$) if
$H(q\mid S_{O})\;\geq\;\omega\bigl(\log|\text{Atom}(S_{O})|\bigr),$ (17)
that is, equivalently,
$\frac{H(q\mid S_{O})}{\log|\text{Atom}(S_{O})|}\;\longrightarrow\;\infty\qquad\text{as }|\text{Atom}(S_{O})|\to\infty.$ (18)
For query classes satisfying the polynomial-depth condition
$Dd(q\mid S_{O})\;\leq\;\mathrm{poly}\bigl(|\text{Atom}(S_{O})|\bigr),$ (19)
(e.g., Datalog, SPARQL, and template-based natural-language queries), conditions (17)–(18) are asymptotically equivalent to
$H(q\mid S_{O})\;\gg\;\log\max\bigl\{|\text{Atom}(S_{O})|,\,Dd(q\mid S_{O})\bigr\},$ (20)
as $|\text{Atom}(S_{O})|\to\infty$.
Lemma 2 (Derivation Incompressibility in the Information-Rich Regime).
Let $S_{O}$ be a knowledge base and $q$ a logical formula such that:
  1. (i)
    $n=Dd(q\mid S_{O})\geq 1$,
  2. (ii)
    $m=|\mathrm{Atom}(S_{O})|$,
  3. (iii)
    $q$ lies in the information-rich regime (Definition 8),
  4. (iv)
    $n=\mathrm{poly}(m)$ (e.g., for bounded-depth reasoning systems).
Then there exists a constant $c_{2}>0$, depending only on the proof system and encoding scheme, such that the conditional Kolmogorov complexity of $q$ given $S_{O}$ satisfies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m).$
By Lemma 1, any derivation of $q$ from $\mathrm{Atom}(S_{O})$ of length $n$ can be encoded in at most
$L_{\text{enc}}=c_{1}\cdot n\log(m+n)+O(\log m)$
bits for some constant $c_{1}>0$. This establishes an upper bound on $K(q\mid\langle\mathrm{Atom}(S_{O})\rangle)$.
Now suppose, for contradiction, that no such lower bound holds. Then for every $\varepsilon>0$, there exist arbitrarily large $m$ and corresponding $q$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)<\varepsilon\cdot n\log(m+n).$
Since $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, so the above implies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)=o\bigl(n\log m\bigr).$
Consequently, the conditional algorithmic entropy satisfies
$H(q\mid S_{O})=K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\cdot\ln 2=o\bigl(n\log m\bigr).$
But because $n\geq 1$ and $n=\mathrm{poly}(m)$, we have $n\log m\geq\log m$, and in fact $n\log m\gg\log m$ as $m\to\infty$. Hence
$H(q\mid S_{O})=o\bigl(n\log m\bigr)\not\gg\log m,$
which contradicts condition (iii)—that $q$ lies in the information-rich regime. Therefore, there must exist a constant $c_{2}>0$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m),$
as claimed. ∎
Throughout this subsection, all asymptotic notation ($O(\cdot)$, $\Theta(\cdot)$, $\omega(\cdot)$, $\gg$, etc.) is taken with respect to
$m\;:=\;|\text{Atom}(S_{O})|\;\longrightarrow\;\infty,$
under the standing assumption that
$n\;:=\;Dd(q\mid S_{O})\;=\;\mathrm{poly}(m).$
Theorem 4 (Derivation Depth as an Information Metric).
Let $S_{O}$ be a knowledge base and let $q$ be a query satisfying conditions (i)(iv) of Lemma 2 (in particular, $q$ lies in the information-rich regime of Definition 8 and $n:=Dd(q\mid S_{O})=\mathrm{poly}(m)$ with $m:=|\text{Atom}(S_{O})|$). Define
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\cdot\ln 2,$ (21)
and, under the classical thermodynamic regime (Assumption 3), define the derivation entropy as
$H_{\mathrm{derive}}(q\mid S_{O})\;:=\;Dd(q\mid S_{O})\cdot\ln 2.$ (22)
Then
@@
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$ (23)
where the implicit constants in the $\Theta(\cdot)$ notation depend only on the choice of universal Turing machine, proof system, and encoding scheme, but not on $S_{O}$ or $q$.
Let $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$. By Lemma 1 and Lemma 2, there exist constants $c_{1},c_{2}>0$ such that
$c_{2}\,n\log(m+n)-O(\log m)\;\leq\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\\ \;\leq\;c_{1}\,n\log(m+n)+O(\log m).$
Multiplying both sides by $\ln 2$ yields
$c_{2}^{\prime}\,n\log(m+n)-O(\log m)\;\leq\;H(q\mid S_{O})\\ \;\leq\;c_{1}^{\prime}\,n\log(m+n)+O(\log m),$
for suitable constants $c_{1}^{\prime},c_{2}^{\prime}>0$ with $c_{i}^{\prime}=c_{i}\ln 2$. Because $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, and the $O(\log m)$ term is asymptotically dominated by $n\log(m+n)$. Hence
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\!\frac{H(q\mid S_{O})}{\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\!\right)\!,$
Rearranging gives
$n\;=\;\Theta\!\left(\frac{H(q\mid S_{O})}{\log(m+n)}\right).$
Finally, substituting $H_{\mathrm{derive}}(q\mid S_{O})=n\cdot\ln 2$ and recalling that $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$, we obtain
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$
which is exactly (23). This completes the proof. ∎
III-B Frequency-weighted storage–computation tradeoff
In this subsection we combine the semantic information measures developed in Section II with the derivation-depth analysis in Section III.A to obtain a quantitative tradeoff between storage and computation under a given query distribution. Intuitively, if a query $q$ will be accessed frequently, it may be advantageous to “precompute and store” part of the information needed to answer $q$; if $q$ is rare, it is better to compute the answer on demand from the base knowledge $S_{O}$. Our main result in this subsection, Theorem 5, formalizes this intuition in terms of semantic information content, derivation entropy, and frequency-weighted amortized cost.
Definition 9 (Access Frequency).
For a query distribution $P_{Q}$ over query set $Q$ and a total of $N$ queries, the access frequency (expected access count) of query $q$ is:
$f_{q}:=N\cdot P_{Q}(q).$ (24)
Thus $f_{q}\in[0,N]$ with $\sum_{q\in Q}f_{q}=N$.
Throughout this subsection we assume Definition 9 has been fixed, and we keep its notation. In particular, we let $P_{Q}$ denote the query distribution over a set $Q$ of admissible queries, and
$f_{q}:=N\cdot P_{Q}(q)$
denote the expected access count of query $q$ over $N$ total queries. [^2] As in Definition 9, $f_{q}$ is a dimensionless quantity: it is a pure count (expected number of accesses) without an intrinsic time unit. When we speak of amortized cost per access, we will divide a total bit-length (storage) or total semantic derivation cost (in nats) by this dimensionless $f_{q}$ to obtain a per-access cost in the corresponding information unit.
We now state the central theorem of this section.
Theorem 5 (Frequency-weighted storage–computation tradeoff).
Let $S_{O}$ be a knowledge base with atomic semantic state set $\text{Atom}(S_{O})$, and let $q$ be a query in the information-rich regime with respect to $S_{O}$ in the sense of Definition 8. Let $P_{Q}$ be a query distribution over a query set $Q$, and define $f_{q}:=N\cdot P_{Q}(q)$ as in Definition 9.
Consider a hybrid strategy $\Pi=(S,C,R)$ where:
@@ III-B Frequency-weighted storage–computation tradeoff
    $S$ is an auxiliary storage structure (e.g., materialized facts, indices, intermediate derivations) that is effectively encodable as a binary string $\text{enc}{S}$ of length $\text{abs}{S}$ bits.
@@
    $C$ and $R$ describe the residual computation (inference) carried out over the combined knowledge base $S_{O}\cup S$.
We impose the following assumptions on $S$ and on inference over $S_{O}\cup S$:
  1. 1.
    (SA1) Information-theoretic measurement of storage. The bit length $|S|$ of the encoding $\text{enc}{S}$ satisfies
    $H(S)\;\leq\;|S|\;\leq\;H(S)+O(1),$ (25)
    where $H(S)$ is the Shannon entropy of $S$ for a suitable prior over admissible storage states. In particular,
    $I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (26)
  2. 2.
    (SA2) Structural regularity of the extended knowledge base. The combined knowledge base $S_{O}\cup S$ admits a well-defined atomic semantic state set $\text{Atom}(S_{O}\cup S)$, a predecessor function satisfying Axioms 1-3, and hence a derivation depth function $Dd(\cdot\mid S_{O}\cup S)$ as in Definition 6. We write
    $m_{S}:=|\text{Atom}(S_{O}\cup S)|,n_{S}:=Dd(q\mid S_{O}\cup S).$ (27)
    We assume $n_{S}=\mathrm{poly}(m_{S})$ (polynomial-depth regime) and that $q$ remains in the information-rich regime with respect to $S_{O}\cup S$ (in the sense of Definition 8, replacing $S_{O}$ by $S_{O}\cup S$). Moreover, the size/depth parameters of $S_{O}\cup S$ are of the same polynomial order as those of $S_{O}$.
@@
    (SA3) Applicability of the derivation-depth bound. The proof-system/encoding conditions (EPS) and (NPC) of Proposition 1 hold not only for the base theory induced by $S_{O}$, but also for the theory induced by $S_{O}\cup S$. Consequently, the asymptotic relation of Theorem 4 applies to inference from $S_{O}\cup S$: there exist constants $c_{1},c_{2}>0$ such that
    $c_{1}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}-O(1)\!\;\leq\;\!H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}+O(1),$ (28)
    where
    $H_{\mathrm{derive}}(q\mid S_{O}\cup S):=Dd(q\mid S_{O}\cup S)\,\lambda$
    is the derivation entropy (Definition 7) with per-step cost $\lambda>0$, and all entropies are measured in nats.
Under these assumptions, define:
  • The semantic information content of $q$ (Definition 8):
    $H_{q}:=H(q\mid S_{O}):=K\big(\text{enc}{q}\mid\text{enc}{\text{Atom}(S_{O})}\big)\,\ln 2.$ (29)
  • The effective derivation cost factor associated with $S$:
    $L_{S}:=\log(m_{S}+n_{S}),$ (30)
    which captures the logarithmic compression advantage available through derivations over $S_{O}\cup S$.
  • The amortized cost per access of answering $q$ using strategy $\Pi$:
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q):=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (31)
Then there exist positive constants $c,C>0$ (depending only on the universal Turing machine, proof system, and encoding conventions, but not on $S_{O}$, $q$, $S$, or $P_{Q}$) such that
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1).$ (32)
In particular, up to additive $O(1)$ and multiplicative constant factors, there exists a strategy $\Pi$:
  1. (1)
    Low-frequency regime ($f_{q}\leq L_{S}$, on-demand computation dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (33)
  2. (2)
    High-frequency regime ($f_{q}>L_{S}$, precomputation/storage dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (34)
The break-even frequency scale is given (up to constant factors) by
$f_{c}\;\asymp\;L_{S}\;=\;\log(m_{S}+n_{S}).$ (35)
Queries with $f_{q}\gg f_{c}$ are asymptotically more cost-effective to support via storage (larger $S$); queries with $f_{q}\ll f_{c}$ are asymptotically more cost-effective to answer via on-demand derivation.
We sketch the main steps.
Step 0: Information-theoretic preliminaries
By (SA1) and the data-processing inequality,
$I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (36)
Using the mutual-information identity
$H(q\mid S_{O})=H(q\mid S_{O},S)+I(q;S\mid S_{O}),$ (37)
we obtain
$H(q\mid S_{O},S)=H_{q}-I(q;S\mid S_{O})\geq H_{q}-|S|.$ (38)
Thus $S$ can reduce the conditional entropy of $q$ (given $S_{O}$) by at most $|S|$ nats.
Step 1: Applying Theorem 4
By (SA2) and (SA3), Theorem 4 applies to $S_{O}\cup S$ and yields
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{L_{S}}+O(1).$ (39)
Using the bounds
$H_{q}-|S|\;\leq\;H(q\mid S_{O}\cup S)\;\leq\;H_{q},$ (40)
we obtain
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (41)
Step 2: Amortized cost
By definition,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (42)
Substituting the previous inequality,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{|S|}{f_{q}}+c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (43)
This decomposition makes explicit the storage-amortization term $|S|/f_{q}$ and the derivation-entropy term $\sim H_{q}/L_{S}$.
Step 3: Extremal strategies
To obtain a clean upper bound in terms of $(H_{q},f_{q},L_{S})$ alone, we compare against two constructive extremes.
(a) Pure computation strategy $\Pi_{\mathrm{comp}}$: take $S=\varnothing$. Then, by Theorem 4 applied to $S_{O}$,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q)=H_{\mathrm{derive}}(q\mid S_{O})\\ \;\asymp\;\frac{H_{q}}{\log\big(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\big)}\\ \;\leq\;C_{0}\,\frac{H_{q}}{L_{S}}+O(1)$ (44)
for some constant $C_{0}>0$, where the last inequality uses that $L_{S}=\log(m_{S}+n_{S})$ is of the same asymptotic order as $\log(|\text{Atom}(S_{O})|+Dd(q\mid S_{O}))$ under the polynomial-depth assumptions.
(b) Pure storage strategy $\Pi_{\mathrm{store}}$: by the coding theorem and the definition of $H_{q}$, there exists a conditional code for the answers to $q$ (given $S_{O}$) of length at most $H_{q}+O(1)$ nats. We can store such a near-optimal representation in $S$, so that $|S|\leq H_{q}+O(1)$, and design $C$ and $R$ so that answering $q$ from $S_{O}\cup S$ requires only constant derivation cost:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)=O(1).$ (45)
Hence
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)=\frac{|S|}{f_{q}}+O(1)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (46)
Since we are interested in achievable costs, the optimal amortized cost
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q):=\inf_{\Pi}\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)$ (47)
is bounded above by the minimum of these two explicit strategies:
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\\ \;\leq\;\min\big\{\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q),\,\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)\big\}\\ \;\leq\;\min\left\{C_{0}\,\frac{H_{q}}{L_{S}},\;\frac{H_{q}}{f_{q}}\right\}+O(1).$ (48)
Equivalently, there exists $C>0$ such that
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\max\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (49)
establishing the claimed bound.
Step 4: Frequency regimes
From the last inequality we obtain two asymptotic regimes:
  • If $f_{q}\leq L_{S}$, then $\frac{H_{q}}{f_{q}}\geq\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (50)
  • If $f_{q}>L_{S}$, then $\frac{H_{q}}{f_{q}}<\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (51)
The transition occurs when $f_{q}$ is on the order of $L_{S}$, i.e., $f_{c}\asymp L_{S}=\log(m_{S}+n_{S})$. ∎
Remark 4 (Physical and engineering interpretation).
Theorem 5 admits a natural physical and engineering interpretation.
(1) Units and dimensions. By construction, $H_{q}$ and $H_{\mathrm{derive}}$ are measured in nats (or bits up to a constant factor), while $|S|$ is measured in bits. The frequency $f_{q}$ is a dimensionless expected count of accesses. Therefore:
  • $|S|/f_{q}$ has the same information dimension as $|S|$ (bits), but represents an average storage cost per access;
  • $H_{\mathrm{derive}}$ is a semantic “work” or “action” measured in nats;
  • the bound in Theorem 5 compares two quantities of the same information-theoretic dimension: per-access storage and per-access semantic computation.
In a physical implementation, both can be converted into expected energy dissipation using Landauer’s principle or related models of irreversible computation.
(2) Two limiting protocols. The proof explicitly constructs two limiting protocols:
@@
    A pure computation protocol, where no auxiliary storage is used ($S=\varnothing$) and the cost is determined entirely by derivation from $S_{O}$; this yields an amortized cost of order $H_{q}/\log(\cdot)$.